# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ E.1. Random LCQP and MI-LCQP instance generation
LCQP instance generation for generalization experiments. In this setting, we only sample different coefficients $c$ for different LCQP instances. We sample other components only once, i.e., $Q$, $A$, $b$, $l$, $u$ and $\circ$ in (2.1), and keep them constant and shared by all instances. We also slightly adjust the distributions from which these components are sampled as described below.
@@
For the generalization experiments, we first generate 25,000 LCQP instances for training, and then take the first 100/500/25,00/5,000/10,000 instances to form the smaller training sets. This ensures that the smaller training sets are subsets of the larger sets. The validation set contains 1,000 instances that are generated separately.
E.2. Details of GNN implementation
@@ E.3. Details of GNN training
(E.1) $L_{\mathcal{G}}(F_{W})=\mathbb{E}_{G\sim\mathcal{G}}\left[\frac{\|F_{W}(G)-\Phi(G)\|_{2}^{2}}{\max(\|\Phi(G)\|,1)^{2}}\right],$
@@
E.4. Full generalization results on LCQP
@@
Figure 7. Training and validation errors when training GNNs with an embedding size of 512 on different numbers of LCQP problem instances to fit $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$.
Figure 7 shows the variations of training and validation errors when training GNNs of an embedding size of 512 on different numbers of LCQP problem instances. We observe similar trends for both prediction tasks, that the generalization gap decreases and the generalization ability improves as more instances are used for training. This result implies the potential of applying trained GNNs to solve QP problems that are unseen during training but are sampled from the same distribution, as long as enough training instances are accessible and the instance distribution is specific enough (in contrast to the generic instances used in experiments of Figure 3 and 4).

### v2

@@ G.1. Random LCQP and MI-LCQP instance generation
LCQP instance generation for generalization experiments. In this setting, we only sample different coefficients $c$ for different LCQP instances. We sample other components only once, i.e., $Q$, AA, $b$, $l$, $u$ and $\circ$ in (2.1), and keep them constant and shared by all instances. We also slightly adjust the distributions from which these components are sampled as described below.
@@
For the generalization experiments, we first generate 25,000 LCQP instances for training, and then take the first 100/500/25,00/5,000/10,000 instances to form the smaller training sets. This ensures that the smaller training sets are subsets of the larger sets. The test set contains 1,000 instances that are generated separately.
Portfolio optimization formulation and instance generation. The portfolio optimization problems are formulated as below.
(G.1) $\min_{x,y}$ $\quad\frac{1}{2}x^{\top}Dx+\frac{1}{2}y^{\top}y-\mu^{\top}x$
s.t. $\quad y=Fx,\quad\mathbf{1}^{\top}x=1,\quad x\geq 0$
Here $x\in\mathbb{R}^{s}$ and $y\in\mathbb{R}^{t}$ are the optimization variables, $D\in\mathbb{R}^{s\times s}$ is a diagonal matrix with non-negative diagonal elements, $F\in\mathbb{R}^{t\times s}$ is the factor modeling matrix. We generate portfolio optimization instances following the scheme below.
  • We set $s=50$ and $t=5$, resulting in LCQP instances with $m=6$ constraints and $n=55$ variables.
  • The diagonal elements of $D$ are independently sampled from uniform distribution: $_{ii}\sim U(0,\sqrt{t}). DD$ is then used to form the matrix $Q=\begin{pmatrix}D&\\ &I_{t}\end{pmatrix}$.
  • The coefficients $\mu$ in the objective function: $\mu_{j}\sim\mathcal{N}(0,1)$.
  • The non-zero elements in the factor modeling matrix $F$: $F_{ij}\sim\mathcal{N}(0,1)$. The coefficient matrix $F$ contains 25 non-zero elements. The positions are sampled randomly.
SVM optimization formulation and instance generation. The support vector machine optimization problems are formulated as below.
(G.2) $\min_{x,t}$ $\quad\frac{1}{2}x^{\top}x+\lambda\mathbf{1}^{\top}t$
s.t. $\quad t\geq\mathrm{diag}(y)Dx+\mathbf{1},\quad t\geq 0$
Here $x\in\mathbb{R}^{s}$ and $t\in\mathbb{R}^{t}$ are the optimization variables, $D\in\mathbb{R}^{t\times s}$ is the data matrix, $y\in\mathbb{R}^{t}$ is the binary label vector, and $\lambda$ is a hyperparameter which we set to $1/2$. We generate SVM optimization instances following the scheme below.
  • We set $s=5$ and $t=50$.
  • The non-zero elements in the data matrix $D$: $D_{ij}\sim\mathcal{N}(-0.1,0.1)$ for $i\leq t/2$; $D_{ij}\sim\mathcal{N}(0.1,0.1)$ otherwise. The coefficient matrix $D$ contains 100 non-zero elements. The positions are sampled randomly.
  • The binary label vector $y$: $y_{i}=-1$ for $i\leq t/2$; $y_{i}=1$ otherwise.
G.2. Details of GNN implementation
@@ G.3. Details of GNN training
(G.3) $L_{\mathcal{G}}(F_{W})=\mathbb{E}_{G\sim\mathcal{G}}\left[\frac{\|F_{W}(G)-\Phi(G)\|_{2}^{2}}{\max(\|\Phi(G)\|,1)^{2}}\right],$
@@
G.4. Generalization results on LCQP
Figure 8 shows the variations of training and test errors when training GNNs of an embedding size of 512 on different numbers of LCQP problem instances. We observe similar trends for both prediction tasks, that the generalization gap decreases and the generalization ability improves as more instances are used for training. This result implies the potential of applying trained GNNs to solve QP problems that are unseen during training but are sampled from the same distribution, as long as enough training instances are accessible and the instance distribution is specific enough (in contrast to the generic instances used in experiments of Figure 3 and 4).
@@ G.4. Generalization results on LCQP
Figure 8. Training and test errors when training GNNs with an embedding size of 512 on different numbers of LCQP instances to fit $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$.
G.5. Details for the Maros-Meszaros test set
To show the fitting ability of GNNs on more realistic QP problems, we train GNNs on the Maros and Meszaros Convex Quadratic Programming Test Problem Set [maros1999repository], which contains 138 quadratic programs that are designed to be challenging. We apply equilibrium scaling to each problem and also scale the objective function so that the $Q$ matrix will not contain too large elements. We collect the optimal solutions and objective values of the test instances using an open-sourced QP solver called PIQP [schwan2023piqp], which is benchmarked to achieve best performances on the Maros Meszaros test set among many other solvers [qpbenchmark2024]. PIQP solves 136 problem instances successfully, which are then used to train four GNNs with embedding size of $64,128,256,512$. The training protocol follows the experiments using synthesized QP instances in Section 5.